# Distribution Grids May Be a Barrier to Residential Electrification

### v3

@@
Priyadarshan Mechanical Engineering, Purdue University, USA Constance Crozier Industrial and Systems Engineering, Georgia Institute of Technology, USA Kyri Baker Civil, Environmental, and Architectural Engineering, University of Colorado Boulder, USA Renewable and Sustainable Energy Institute, University of Colorado Boulder, USA Kevin J. Kircher Mechanical Engineering, Purdue University, USA Electrical and Computer Engineering, Purdue University, USA Lead contact Correspondence: kircher@purdue.edu
@@
Replacing fossil-fueled appliances and vehicles with electric alternatives can reduce greenhouse gas emissions and air pollution in many settings. However, residential electrification can also raise electricity demand beyond the safe limits of electrical infrastructure. This can increase the risk of blackouts or require grid reinforcement that is often slow and expensive. Here, we estimate the physical and economic impacts on distribution grids of electrifying all housing and personal vehicles in each county of the lower 48 United States. We find that space heating is the main driver of grid impacts, with the coldest regions seeing demand peaks up to five times higher than today’s peaks. Accommodating electrification of all housing and personal vehicles is estimated to require 600 GW of distribution grid reinforcement nationally, at a cost of \$350 to \$790 billion, or \$2,800 to \$6,400 per household (95% confidence intervals). However, demand-side management could eliminate three-quarters of grid reinforcement costs.
@@ INTRODUCTION
Fossil-fueled driving, space heating, and water heating account for about one-third of United States greenhouse gas emissions 1. Replacing fossil-fueled vehicles and heaters with electric alternatives can significantly reduce these emissions today 2, and could essentially eliminate them if the United States achieves its goal of 100% carbon-free electricity by 2035 3. For these reasons, the Inflation Reduction Act of 2022 incentivizes adoption of electric vehicles and heat pumps (efficient electric heating and cooling machines) through tax credits and point-of-sale rebates 4, 5, 6. State and local governments have enacted complementary incentives, including some bans on fossil-fueled vehicles or heaters 7. Due in part to these policies, adoption of electric vehicles and heat pumps in the United States is accelerating 8.
@@
While electrification of fossil-fueled activities can deeply reduce emissions, it could also have serious impacts on the power grid. Today, electricity demand in most United States power systems peaks in the hottest hours of the year, driven mainly by air conditioning 9. Electrifying space heating can heighten annual demand peaks and shift them from summer to winter 10, 11, 12, 13, 14. Electric vehicle adoption is likely to further exacerbate demand peaks 15, 16, 17. Heightened demand peaks from electrification could cause power quality issues and make power outages more likely 18, 15. Mitigating these risks often requires replacing power lines and transformers, which are typically sized to accommodate power flows in near-worst-case demand forecast scenarios. Transformers in particular are already in short supply, with United States utilities experiencing lead times of up to two years and prices that have increased by 400 to 900% in the last three years 19. Utilities pass grid reinforcement costs on to ratepayers in the form of increased electricity prices 20, 21, which weaken economic incentives for further electrification.
@@
Past research on the grid impacts of electrification focused mainly on high-voltage transmission systems. At the nationwide scale, the National Renewable Energy Laboratory’s Electrification Futures Study 22, 10, 23 investigated how electrification might alter load profiles and transmission needs. In their high-electrification scenario (about 90% adoption of electric vehicles and 70% of heat pumps), peak electricity demand rose by 38%. A nationwide study from the Electric Power Research Institute 24 estimated that 75% adoption of electric vehicles and 50% of heat pumps (by floor area) would raise peak demand by 24% to 52%. Wilson et al. 2 used the ResStock tool 25 to investigate nationwide cost and emission impacts under various heat pump adoption scenarios. While Wilson et al. 2 did not explicitly consider grid impacts, they stressed the need for future work to understand how evolving load profiles should shape infrastructure investments. At the regional scale, an American Council for an Energy-Efficient Economy 26 study found that 30% adoption of EVs and 100% of heat pumps could increase peak demand in New England by 235%. White et al. 27 estimated that electrifying 100% of space heating would require 25% expansion of Texas grid capacity. Zhang et al. 28 estimated that 10% adoption of electric vehicles with uncoordinated charging would increase Midwest peak demand by 10%.
@@
In parallel with high-voltage transmission grid research at the nationwide and regional scales, several studies investigated medium-voltage distribution grid impacts in California specifically. Elmallah et al. 29 used the Distribution Deferral Opportunity Report, which the California Public Utilities Commission requires utilities to file each year, to analyze electrification impacts within the Pacific Gas & Electric service territory. Elmallah et al. estimated that distribution grid reinforcement to accommodate new load from electrification, including up to 100% adoption of electric vehicles and up to 50% adoption of heat pumps and water heaters, could cost \$1 to \$10 billion (about \$200 to \$2,000 per customer). Li et al. 30 used the same dataset to study 100% adoption of electric vehicles (but not heat pumps or water heaters) in California, estimating that 67% of distribution feeders could need reinforcement, at a total cost of \$6 to \$20 billion (about \$300 to \$900 per customer). Several other state-specific studies investigated the impacts of electrification on greenhouse gas emissions without considering impacts on power grids 14, 31, 32.
@@
We estimate distribution grid reinforcement requirements through bottom-up simulation. We estimate today’s distribution grid capacity in each county by simulating all the appliances and vehicles in 1,000 representative housing units during peak heating and cooling weeks, rescaling the overall peak aggregate demand to reflect the true number of housing units, and adding a safety margin to represent headroom in today’s capacity. We repeat the same process for an all-electric future. The estimated grid reinforcement requirement is the difference between the all-electric future grid capacity and today’s grid capacity. To simulate buildings and vehicles, we use the open-source EDGIE (Emulating the Distribution Grid Impacts of Electrification) toolbox 33. As illustrated in Fig. 1, EDGIE models heat pumps, electric vehicles, water heaters, and building thermal dynamics. It uses real data on weather 34, indoor temperature setpoints 35, domestic hot water use 36, and miscellaneous electrical loads 37. EDGIE implements empirically validated, multi-physics modeling at hourly or sub-hourly time resolution and county-level spatial resolution. We tune building, appliance, and vehicle parameters to county-specific data. We make all data publicly available to facilitate reproduction or extension of the work.
@@ RESULTS
We assess the physical impacts of home and vehicle electrification on distribution grids through bottom-up modeling. In each county, we simulate all the residential appliances and personal vehicles from 1,000 representative households, then rescale the aggregate load profile to reflect the true number of households. Simulated appliance and vehicle properties, housing sizes and types, and levels of insulation and air sealing all vary both across and within counties. We fit physical and behavioral parameters associated with buildings, thermal equipment, and vehicles to county- and climate-specific data, using real data wherever possible. The Experimental Procedures section and the Supplementary Information document provide more details on models and data sources.
We simulate peak heating and cooling weeks under two scenarios: (1) a business-as-usual scenario that reflects today’s building stock, appliances, and vehicles; and (2) an all-electric scenario that reflects a possible future with complete electrification of personal vehicles and residential space and water heating. In the all-electric scenario, each home has a central air-to-air heat pump with backup resistance heat. We compute design heating and cooling loads for each home under the 99% heating and 1% cooling design temperatures from the American Society of Heating, Refrigeration, and Air-Conditioning Engineers 2017 Handbook of Fundamentals and the Air Conditioning Contractors of America Manual J Design Conditions 8th Edition. Each central heat pump is sized to meet the larger of the home’s design heating load and design cooling load, up to a cap based on today’s equipment sizes. We model the largest available central heat pump as having five tons (17.6 kW) of nameplate cooling capacity. If the design heating or cooling load exceeds the maximum capacity of the largest available central heat pump under design conditions, we add a mini-split heat pump with one to three tons (3.5 to 10.6 kW) of nameplate cooling capacity. Resistance backup meets any heat demand in excess of the combined capacities of the central heat pump and the mini-split heat pump. In the two northernmost climate zones, we simulate adoption of heat pumps whose coefficients of performance meet the United States Department of Energy’s Cold Climate Heat Pump Challenge specifications 38. The modeled heat pump capacities and coefficients of performance vary with outdoor temperatures.
@@
We estimate the distribution grid reinforcement requirement as the difference between the all-electric and business-as-usual distribution grid capacities. We estimate the distribution grid capacities in each county and scenario by simulating the appliances and vehicles for 1,000 households, rescaling the aggregate demand to the true number of households, taking the 99th percentile, then adding a safety margin to reflect distribution grid headroom. We base headroom estimates on a hosting capacity analysis for seven utilities in New York State 39, which found that today’s headroom varies by utility from 15% to 36%. We use this headroom range for the business-as-usual scenario, randomizing over counties. For the all-electric scenario, we assume all counties have 20% headroom. The discussion section analyzes the sensitivity of grid reinforcement requirements to headroom assumptions in the business-as-usual and all-electric scenarios.
@@
We draw on three studies 29, 40, 41 to determine a plausible price range for typical distribution grid reinforcement, which may include installing power lines, transformers, switchgear, capacity banks, or other infrastructure. Based on Pacific Gas & Electric data, Elmallah et al. 29 estimated prices of 860 to 36,500 \$/kW, with the highest prices for the smallest projects. (We inflation-adjust all price data to 2024 dollars.) Zhang et al. 40 used Energy Information Administration data 42 to derive state-wise prices and a national average of 1,664 \$/kW, including a variety of administrative costs. Rauschkolb et al. 41 analyzed data from the Federal Energy Regulatory Commission’s Form 1 43, where utilities can report their capital, operation, and maintenance costs related to grid reinforcement. By contrast to prior studies, Raushkolb et al. separated the regular costs of sustaining an existing distribution grid from the incremental costs of expanding grid capacity to accommodate load growth. Rauschkolb et al. developed a linear model that accounts for both one-time capital costs from expanding grid capacity, and for increased annual costs from operating and maintaining a larger grid.
Here, we modify Rauschkolb et al.’s approach to estimate a representative price range for distribution grid reinforcement. We assume grid capacities increase linearly from the business-as-usual capacities in 2025 to the all-electric capacities in 2050. We map future costs into net present costs using discounted cash flow analysis. Similar to Rauschkolb et al.’s approach, our modified approach accounts for both one-time capital costs of reinforcing infrastructure, and increased annual costs from operating and maintaining a larger grid. The Experimental Procedures section contains the details of our modifications, which yield a mean grid reinforcement price of 960 \$/kW and a 95% confidence interval of 587 to 1,331 \$/kW.
@@
The grid reinforcement cost estimates described above represent incremental spending, above and beyond the business-as-usual spending required to sustain existing distribution grids. The nationwide grid reinforcement cost estimate over 25 years in the central price scenario is \$569 billion. For comparison, we analyzed distribution grid spending data from the Federal Energy Regulatory Commission’s Form 1 43 from 1995 to 2019. Over that 25-year period, the total inflation-adjusted spending on distribution grid infrastructure – including all capital, operation, and maintenance costs associated with both sustaining existing grids and expanding capacity to accommodate load growth from residential, commercial, and industrial customers – was \$975 billion. Relative to this cost, the estimated incremental grid reinforcement cost of \$569 billion to accommodate home and vehicle electrification is a 58% increase. While this is not an exact comparison, it gives an approximate sense of scale for the incremental distribution grid spending required to accommodate residential electrification.
Aggregated over the lower 48 United States, distribution grid reinforcement to accommodate 100% electrification of homes and private vehicles over 25 years is estimated to cost an additional \$350 to \$790 billion (\$110 to \$250 per household per year), above and beyond business-as-usual spending. For context, United States utilities spent \$975 billion on distribution grid infrastructure in the 25 years from 1995 to 2019 (\$390 per customer per year) 43. Residential energy bills cost Americans about \$230 billion annually (\$1,800 per household per year) 44. The Inflation Reduction Act of 2022 allocated about \$370 billion over ten years (\$290 per household per year) to spending on energy and climate 45. The United States annual military budget was about \$820 billion (\$6,500 per household per year) in 2023 46.
@@
To assess the cost reduction potential of smart electrification, we simulate three demand-side management strategies. First, thermal demand from buildings could decrease due to better insulation and air sealing, due to shifts from detached housing to less energy-intensive attached housing, and/or due to shifts to smaller housing. We simulate thermal demand reduction by increasing the effective thermal resistance (which models the combined effects of wall and roof insulation, window quality, outdoor air infiltration rates, and the surface area exposed to the outdoor air) by 25%, consistent with the ‘basic enclosure package’ from Maxim and Grubert 47. Second, the air-source heat pumps simulated above could shift to ground-source heat pumps, which use about one-third less electricity per unit of heat output and require less backup resistance heat. Third, software could reduce peak aggregate demand by coordinating the operation of water heaters, electric vehicles, and heat pumps. For example, heat pumps could preheat homes in anticipation of aggregate demand peaks, or electric vehicles could interleave their charging with heat pump operation. (We simulate device coordination through convex optimization with perfect information in one representative location per state, as discussed in the Supplementary Information document.) The demand-side management strategies simulated here are far from exhaustive. We leave for future work a range of measures, such as solar photovoltaics, bidirectional electric vehicle charging, home batteries, and thermal storage, that could further reduce grid reinforcement requirements.
@@
Fig. 6 shows nationwide grid reinforcement cost reductions from the three individual strategies and combinations thereof. These estimates use the mean grid reinforcement price of 960 \$/kW, which gives a nationwide cost of \$569 billion (\$180 per household per year) without demand-side management. Individually (top set of bars), the three strategies reduce costs by 19%, 28% and 43%, respectively. Perhaps surprisingly, device coordination – a strategy based mainly on software, sensing, and communication, rather than equipment upgrades – reduces costs more than switching to ground-source heat pumps, a significantly more hardware-intensive and expensive strategy. Cost savings from combined strategies (bottom set of bars) are not strictly additive. For example, the combination of less thermal demand from buildings (19% individually) and ground-source heat pumps (28%) reduces costs by 40%. Combining all three demand-side management strategies reduces nationwide grid reinforcement costs by 71%, to a total of \$165 billion (\$50 per household per year).

### v4

@@
Priyadarshan Mechanical Engineering, Purdue University, West Lafayette, IN, USA Constance Crozier Industrial and Systems Engineering, Georgia Institute of Technology, Atlanta, GA, USA Kyri Baker Civil, Environmental, and Architectural Engineering, University of Colorado Boulder, Boulder, CO, USA Renewable and Sustainable Energy Institute, University of Colorado Boulder, Boulder, CO, USA Kevin J. Kircher Mechanical Engineering, Purdue University, West Lafayette, IN, USA Electrical and Computer Engineering, Purdue University, West Lafayette, IN, USA Lead contact Correspondence: kircher@purdue.edu
@@
Replacing fossil-fueled appliances and vehicles with electric alternatives can reduce greenhouse gas emissions and air pollution in many settings. However, electrification can also raise electricity demand beyond the safe limits of electrical infrastructure. This can increase the risk of blackouts or require grid reinforcement that is often slow and expensive. Here we estimate the physical and economic impacts on distribution grids of electrifying all housing and personal vehicles in each county of the lower 48 United States. We find that space heating is the main driver of grid impacts, with the coldest regions seeing demand peaks up to five times higher than today’s peaks. Accommodating electrification of all housing and personal vehicles is estimated to require 600 GW of distribution grid reinforcement nationally, at a cost of \$350 to \$790 billion, or \$2,800 to \$6,400 per household (95% confidence intervals). However, demand-side management could eliminate over two-thirds of grid reinforcement costs.
@@ INTRODUCTION
Fossil-fueled driving, space heating, and water heating account for about half of United States greenhouse gas emissions 1. Replacing fossil-fueled vehicles and heaters with electric alternatives can significantly reduce these emissions today 2, and could essentially eliminate them in a future with 100% carbon-free electricity. For these reasons, state and local governments have encouraged electrification through a range of policies, including subsidies for electric vehicles and appliances, as well as some bans on new fossil-fueled vehicles or appliances 3. Due in part to these policies, adoption of electric vehicles and appliances in the United States is accelerating 4.
@@
While electrification of fossil-fueled activities can deeply reduce emissions, it could also have serious impacts on the power grid. Today, electricity demand in most United States power systems peaks in the hottest hours of the year, driven mainly by air conditioning 5. Electrifying space heating can heighten annual demand peaks and shift them from summer to winter 6, 7, 8, 9, 10. Electric vehicle adoption is likely to further exacerbate demand peaks 11, 12, 13. Heightened demand peaks from electrification could cause power quality issues and make power outages more likely 14, 11. Mitigating these risks often requires replacing power lines and transformers, which are typically sized to accommodate power flows in near-worst-case demand forecast scenarios. Transformers in particular are already in short supply, with United States utilities experiencing lead times of up to two years and prices that have increased by 400 to 900% in the last three years 15. Utilities pass grid reinforcement costs on to ratepayers in the form of increased electricity prices 16, 17, which weaken economic incentives for further electrification. While increased electricity sales from electrification of heating and other end uses may cover some of these costs, and can even put downward pressure on retail electricity prices18, the overall economic implications of distribution grid reinforcement under deep electrification remain understudied.
@@
Past research on the grid impacts of electrification focused mainly on high-voltage transmission systems. At the nationwide scale, the National Renewable Energy Laboratory’s Electrification Futures Study 19, 6, 20 investigated how electrification might alter load profiles and transmission needs. In their high-electrification scenario (about 90% adoption of electric vehicles and 70% of heat pumps), peak electricity demand rose by 38%. A nationwide study from the Electric Power Research Institute 21 estimated that 75% adoption of electric vehicles and 50% of heat pumps (by floor area) would raise peak demand by 24% to 52%. Wilson et al. 2 used the ResStock tool 22 to investigate nationwide cost and emission impacts under various heat pump adoption scenarios. While Wilson et al. 2 did not explicitly consider grid impacts, they stressed the need for future work to understand how evolving load profiles should shape infrastructure investments. At the regional scale, an American Council for an Energy-Efficient Economy 23 study found that 30% adoption of EVs and 100% of heat pumps could increase peak demand in New England by 235%. White et al. 24 estimated that electrifying 100% of space heating would require 25% expansion of Texas grid capacity. Zhang et al. 25 estimated that 10% adoption of electric vehicles with uncoordinated charging would increase Midwest peak demand by 10%.
@@
In parallel with high-voltage transmission grid research at the nationwide and regional scales, several studies investigated medium-voltage distribution grid impacts in specific states or for vehicle electrification only. Elmallah et al. 26 used the Distribution Deferral Opportunity Report, which the California Public Utilities Commission requires utilities to file each year, to analyze electrification impacts within the Pacific Gas & Electric service territory. Elmallah et al. estimated that distribution grid reinforcement to accommodate new load from electrification, including up to 100% adoption of electric vehicles and up to 50% adoption of heat pumps and water heaters, could cost \$1 to \$10 billion (about \$200 to \$2,000 per household). Li et al. 27 used the same dataset to study 100% adoption of electric vehicles (but not heat pumps or water heaters) in California, estimating that 67% of distribution feeders could need reinforcement, at a total cost of \$6 to \$20 billion (about \$300 to \$900 per household). Mai et al. estimated statewide distribution grid reinforcement costs in New York with up to 100% electric vehicle adoption at \$1.4 to \$26.8 billion (about \$200 to \$3,500 per household) in their lowest- to highest-impact scenarios 28. Cutter et al. estimated that 100% electric vehicle adoption across the United States could require up to \$200 billion in distribution grid reinforcement (about \$1,500 per household) 29.Several other state-specific studies investigated the impacts of electrification on greenhouse gas emissions without considering impacts on power grids 10, 30, 31.
@@
We estimate distribution grid reinforcement requirements through bottom-up simulation. We estimate today’s distribution grid capacity in each county by simulating all the appliances and vehicles in 1,000 representative housing units during peak heating and cooling weeks, rescaling the overall peak aggregate demand to reflect the true number of housing units, and adding a safety margin to represent headroom in today’s capacity. We repeat the same process for an all-electric future. The estimated grid reinforcement requirement is the difference between the all-electric future grid capacity and today’s grid capacity. To simulate buildings and vehicles, we use the open-source EDGIE (Emulating the Distribution Grid Impacts of Electrification) toolbox 32. As illustrated in Fig. 1, EDGIE models heat pumps, electric vehicles, water heaters, and building thermal dynamics. It uses real data on weather 33, indoor temperature setpoints 34, domestic hot water use 35, and miscellaneous electrical loads 36. EDGIE implements empirically validated, multi-physics modeling at hourly or sub-hourly time resolution and county-level spatial resolution. We tune building, appliance, and vehicle parameters to county-specific data. We make all data publicly available to facilitate reproduction or extension of the work.
@@ RESULTS
We assess the physical impacts of home and vehicle electrification on distribution grids through bottom-up modeling. In each county, we simulate all the residential appliances and personal vehicles from 1,000 representative households, then rescale the aggregate load profile to reflect the true number of households. Simulated appliance and vehicle properties, housing sizes and types, and levels of insulation and air sealing all vary both across and within counties. We fit physical and behavioral parameters associated with buildings, thermal equipment, and vehicles to county- and climate-specific data, using real data wherever possible. The Methods section and the Supplemental Notes provide more details on models and data sources (Fig. S1-S7).
We simulate peak heating and cooling weeks under two scenarios: (1) a business-as-usual scenario that reflects today’s building stock, appliances, and vehicles; and (2) an all-electric scenario that reflects a possible future with complete electrification of personal vehicles and residential space and water heating. In the all-electric scenario, each home has a central air-to-air heat pump with backup resistance heat. We compute design heating and cooling loads for each home under the 99% heating and 1% cooling design temperatures from the American Society of Heating, Refrigeration, and Air-Conditioning Engineers 2017 Handbook of Fundamentals and the Air Conditioning Contractors of America Manual J Design Conditions 8th Edition. Each central heat pump is sized to meet the larger of the home’s design heating load and design cooling load, up to a cap based on today’s equipment sizes. We model the largest available central heat pump as having five tons (17.6 kW) of nameplate cooling capacity. If the design heating or cooling load exceeds the maximum capacity of the largest available central heat pump under design conditions, we add a mini-split heat pump with one to three tons (3.5 to 10.6 kW) of nameplate cooling capacity. Resistance backup meets any heat demand in excess of the combined capacities of the central heat pump and the mini-split heat pump. In the two northernmost climate zones, we simulate adoption of heat pumps whose coefficients of performance meet the United States Department of Energy’s Cold Climate Heat Pump Challenge specifications 37. The modeled heat pump capacities and coefficients of performance vary with outdoor temperatures.
@@
We estimate the distribution grid reinforcement requirement as the difference between the all-electric and business-as-usual distribution grid capacities. We estimate the distribution grid capacities in each county and scenario by simulating the appliances and vehicles for 1,000 households, rescaling the aggregate demand to the true number of households, taking the 99th percentile, then adding a safety margin to reflect distribution grid headroom. We base headroom estimates on a hosting capacity analysis for seven utilities in New York State 38, which found that today’s headroom varies by utility from 15% to 36%. We use this headroom range for the business-as-usual scenario, randomizing over counties. For the all-electric scenario, we assume all counties have 20% headroom. The discussion section analyzes the sensitivity of grid reinforcement requirements to headroom assumptions in the business-as-usual and all-electric scenarios.
@@
We draw on three studies 26, 39, 40 to determine a plausible price range for typical distribution grid reinforcement, which may include installing power lines, transformers, switchgear, capacity banks, or other infrastructure. Based on Pacific Gas & Electric data, Elmallah et al. 26 estimated prices of 860 to 36,500 \$/kW, with the highest prices for the smallest projects. (We inflation-adjust all price data to 2024 dollars.) Zhang et al. 39 used Energy Information Administration data 41 to derive state-wise prices and a national average of 1,664 \$/kW, including a variety of administrative costs. Rauschkolb et al. 40 analyzed data from the Federal Energy Regulatory Commission’s Form 1 42, where utilities can report their capital, operation, and maintenance costs related to grid reinforcement. By contrast to prior studies, Raushkolb et al. separated the regular costs of sustaining an existing distribution grid from the incremental costs of expanding grid capacity to accommodate load growth. Rauschkolb et al. developed a linear model that accounts for both one-time capital costs from expanding grid capacity, and for increased annual costs from operating and maintaining a larger grid.
Here, we modify Rauschkolb et al.’s approach to estimate a representative price range for distribution grid reinforcement. We assume grid capacities increase linearly from the business-as-usual capacities in 2025 to the all-electric capacities in 2050. We map future costs into net present costs using discounted cash flow analysis. Similar to Rauschkolb et al.’s approach, our modified approach accounts for both one-time capital costs of reinforcing infrastructure, and increased annual costs from operating and maintaining a larger grid. This approach captures distribution grid reinforcement costs incurred at primary components, such as substation transformers and high-capacity feeders, as well as secondary components, such as final line transformers and lower-capacity lines. Our approach implicitly assumes that reinforcement requirements for both primary and secondary components scale with the increase in system-wide peak demand. This assumption may lead to overestimating the costs of reinforcing secondary components, which experience demand peaks that may not coincide with system-wide peaks. However, this overestimation is likely small, as the \$/kW price of reinforcing secondary components is about two orders of magnitude lower than the \$/kW price of reinforcing primary components43. We do not model electrical infrastructure costs typically borne by individual customers, such as replacing service drops, circuit breaker panels, or wiring within buildings 44. The Methods section contains the details of our modifications, which yield a mean grid reinforcement price of 960 \$/kW and a 95% confidence interval of 587 to 1,331 \$/kW.
The cost of expanding distribution capacity can vary depending on whether it involves extending infrastructure to new construction vs. retrofitting existing infrastructure to serve higher demand from the same customers. For instance, replacing an aging transformer with a higher-capacity transformer during routine maintenance, or upsizing a line during undergrounding for risk mitigation, may add significant capacity at a comparatively low incremental cost. The Raushkolb et al. model used here was fit to cost data from a period where distribution grid expansion was driven mainly by new construction, rather than retrofitting existing infrastructure. As grid reinforcement to accommodate electrification will likely involve substantial retrofitting of existing infrastructure, our cost estimates may be somewhat high.
@@
The grid reinforcement cost estimates described above represent incremental spending, above and beyond the business-as-usual spending required to sustain existing distribution grids. The nationwide grid reinforcement cost estimate over 25 years in the central price scenario is \$570 billion. For comparison, we analyzed distribution grid spending data from the Federal Energy Regulatory Commission’s Form 1 42 from 1995 to 2019. Over that 25-year period, the total inflation-adjusted spending on distribution grid infrastructure – including all capital, operation, and maintenance costs associated with both sustaining existing grids and expanding capacity to accommodate load growth from residential, commercial, and industrial customers – was \$975 billion. Relative to this cost, the estimated incremental grid reinforcement cost of \$570 billion to accommodate home and vehicle electrification is a 58% increase. While this is not an exact comparison, it gives an approximate sense of scale for the incremental distribution grid spending required to accommodate residential electrification.
Aggregated over the lower 48 United States, distribution grid reinforcement to accommodate 100% electrification of homes and private vehicles over 25 years is estimated to cost an additional \$350 to \$790 billion (\$110 to \$250 per household per year), above and beyond business-as-usual spending. For context, United States utilities spent \$975 billion on distribution grid infrastructure in the 25 years from 1995 to 2019 (\$390 per customer per year) 42. Residential energy bills cost Americans about \$230 billion annually (\$1,800 per household per year) 45. The Inflation Reduction Act of 2022 allocated about \$370 billion over ten years (\$290 per household per year) to spending on energy and climate 46. The United States annual military budget was about \$820 billion (\$6,500 per household per year) in 2023 47.
@@
To assess the cost reduction potential of smart electrification, we simulate three demand-side management strategies. First, thermal demand from buildings could decrease due to better insulation and air sealing, due to shifts from detached housing to less energy-intensive attached housing, and/or due to shifts to smaller housing. We simulate thermal demand reduction by increasing the effective thermal resistance (which models the combined effects of wall and roof insulation, window quality, outdoor air infiltration rates, and the surface area exposed to the outdoor air) by 25%, consistent with the ‘basic enclosure package’ from Maxim and Grubert 48. Second, the air-source heat pumps simulated above could shift to ground-source heat pumps, which use about one-third less electricity per unit of heat output and require less backup resistance heat. Third, software could reduce peak aggregate demand by coordinating the operation of water heaters, electric vehicles, and heat pumps. For example, heat pumps could preheat homes in anticipation of aggregate demand peaks, or electric vehicles could interleave their charging with heat pump operation. (We simulate device coordination through convex optimization with perfect information in one representative location per state, as discussed in the Supplemental Notes.) The demand-side management strategies simulated here are far from exhaustive. We leave for future work a range of measures, such as solar photovoltaics, bidirectional electric vehicle charging, home batteries, and thermal storage, that could further reduce grid reinforcement requirements.
@@
Fig. 6 shows nationwide grid reinforcement cost reductions from the three individual strategies and combinations thereof. These estimates use the mean grid reinforcement price of 960 \$/kW, which gives a nationwide cost of \$570 billion (\$180 per household per year) without demand-side management. Individually (top set of bars), the three strategies reduce costs by 19%, 28% and 43%, respectively. Perhaps surprisingly, device coordination – a strategy based mainly on software, sensing, and communication, rather than equipment upgrades – reduces costs more than switching to ground-source heat pumps, a significantly more hardware-intensive and expensive strategy. Cost savings from combined strategies (bottom set of bars) are not strictly additive. For example, the combination of less thermal demand from buildings (19% individually) and ground-source heat pumps (28%) reduces costs by 40%. Combining all three demand-side management strategies reduces nationwide grid reinforcement costs by 71%, to a total of \$165 billion (\$50 per household per year).